Homework: 9-3: Two Dependent Samples (Matched Pairs) Save Score: 0 of 20 pts 3 of 3 (2 complete) HW Score: 51.28%, 32.31 of 6Identify the test statistic.

Identify the​ P-value.

What is the conclusion based on the hypothesis​test?
Answer #1

The null and alternate hypothesis are:

H0: \mu_{d}=0

Ha: \mu_{d}>0








TOTAL
President176184178176185170
Opponent162181171172183186
d143742-1614
d2196949164256530

\overline{x}_{d}=\frac{\sum x_{d}}{n}=\frac{14}{6}=2.3333

S_{d}=\sqrt{\frac{\sum x_{d}^2-n\overline{x}_{d}^2}{n-1}} =\sqrt{\frac{530-6(2.3333)^2}{6-1}}=9.9733

The test statistic is given by:

t=\frac{\overline{x}_{d}}{\frac{S_{d}}{\sqrt{n}}}=\frac{2.3333}{\frac{9.9733}{\sqrt{6}}}=0.57

Since this is a right-tailed test, so the p-value is given by:

P(t_{n-1}>0.57)=P(t_{5}>0.57)=0.296667

Since p-value is greater than 0.05, so we do not have sufficient evidence to reject the null hypothesis
H0.

Thus we cannot say that \mu_{d}>0 i.e. difference in greater than 0 cm.


answered by: ANURANJAN SARSAM